### v2

# thanks: Corresponding author

Abstract
Type Ia supernovae (SNe Ia) serve as crucial cosmological distance indicators because of their empirical consistency in peak luminosity and characteristic light curve decline rates. These properties facilitate them to be standardized candles for the determination of the Hubble constant ($H_{0}$) within late-time universe cosmology. Nevertheless, a statistically significant difference persists between $H_{0}$ values derived from early and late-time measurements, a phenomenon known as the Hubble tension. Furthermore, recent observations have identified a subset of over-luminous SNe Ia, characterized by peak luminosities exceeding the nominal range and faster decline rates. These discoveries raise questions regarding the reliability of SNe Ia as standard candles for measuring cosmological distances. In this article, we present the Bayesian analysis of eight over-luminous SNe Ia and show that they yield a lower $H_{0}$ estimate, exhibiting closer concordance with $H_{0}$ estimates derived from early-universe data. This investigation potentially represents a step toward addressing the Hubble tension.
1 Introduction
The stability of a main-sequence star is achieved through dynamical equilibrium between the inward gravitational force and the outward radiation pressure generated by nuclear fusion processes within the stellar core. As a star burns out of its nuclear fuel, the radiation pressure diminishes, leading to gravitational contraction. The dynamics of this contraction are significantly influenced by quantum mechanical effects, specifically the onset of electron degeneracy. For main-sequence stars with initial masses less than approximately $10\pm 2M_{\odot}$ Lauffer et al. (2018), the electron degeneracy pressure can effectively counterbalance the gravitational pull, preventing further collapse. The resulting stellar remnant, composed of degenerate electron matter, is a white dwarf (WD). The maximum mass that a non-rotating non-magnetized carbon-oxygen WD can hold is approximately $1.4M_{\odot}$, which is famously known as the Chandrasekhar mass limit Chandrasekhar (1931, 1935).
If a WD accretes matter that exceeds this maximum mass, it becomes dynamically unstable. The temperature at the core of the WD increases to initiate carbon and oxygen fusion, leading to the rapid synthesis of heavier elements. This thermonuclear runaway results in a powerful and luminous explosion, classified as type Ia supernova (SN Ia). These explosions typically release approximately $10^{44}\rm\,J$ energy Liu et al. (2023). They are characterized by their extreme luminosity, rendering them observable over vast cosmological distances. Spectroscopically, SNe Ia are identified by the absence of hydrogen and helium emission lines, but exhibit a distinct silicon absorption line at approximately 6150 Å. While each individual SN Ia exhibits variations in its light curve, the dispersion in their peak luminosities is minimal, attributed to their progenitors reaching near the Chandrasekhar mass limit.
SNe Ia exhibit highly consistent light curves, characterized by predictable temporal variations in their peak luminosities, making them valuable cosmological distance indicators, often referred to as ‘standard candles.’ Distance estimations are carried by comparing the absolute and apparent magnitudes, a concept first proposed in Supernova Cosmology Project et al. (1998). A well-established empirical relationship, known as the Phillips relation Phillips (1993), correlates with a higher peak luminosity with slower luminosity decline rates in the light curve. This inherent property has led to extensive utilization of SNe Ia in the determination of the Hubble constant ($H_{0}$) within the local universe. $H_{0}$ is a fundamental cosmological parameter, providing insights into the expansion rate and age of the universe. Using several SNe Ia data along with measurements from cephid variable stars, the SH0ES collaboration Riess et al. (2022), obtained $H_{0}=73.04\pm 1.04\rm\,km\,s^{-1}\,Mpc^{-1}$. However, the analysis of cosmic microwave background (CMB) data obtained by the Planck satellite, within the framework of the $\Lambda$ cold dark matter ($\Lambda$CDM) cosmological model, yields $H_{0}=67.4\pm 0.5\rm\,km\,s^{-1}\,Mpc^{-1}$ Planck Collaboration et al. (2020). This discrepancy between $H_{0}$ measurements derived from early- and late-time universe observations is commonly referred to as the Hubble tension. A comprehensive review of the various methodologies proposed to resolve this tension can be found in Di Valentino et al. (2021).
Over the last couple of decades, observations have revealed a population of over-luminous SNe Ia, characterized by exceptionally high peak luminosities and faster decay rate of the light curve than expected Howell et al. (2006); Hicken et al. (2007); Scalzo et al. (2010); Scalzo and Supernova Factory (2012); Yamanaka et al. (2009); Silverman et al. (2011); Taubenberger et al. (2011); Parrent et al. (2016); Chakradhari et al. (2014); Cao et al. (2016); Dimitriadis et al. (2022); Hsiao et al. (2020). It was argued that they potentially originate from WDs with masses exceeding the standard Chandrasekhar mass limit Miller (2010). Several theoretical mechanisms have been proposed to explain the formation of these super-Chandrasekhar mass WDs. These mechanisms include the presence of strong magnetic fields exceeding the Schwinger limit, which enhance the magnetic pressure, thereby allowing for increased mass accumulation Das and Mukhopadhyay (2013a, b). Another hypothesis involves high angular momentum, where rapid spin generates a centrifugal force that expands the WD, enabling further mass accretion while maintaining hydrostatic equilibrium Fink et al. (2018). Additionally, modified gravity theories, which effectively alter the Poisson equation, have been proposed as a means of producing WDs with masses significantly exceeding $1.4M_{\odot}$ Kalita and Mukhopadhyay (2018, 2021); Kalita and Sarmah (2022). Furthermore, the effects of noncommutative geometry, which become significant at length scales comparable to the electron Compton wavelength, can modify the equation of state of degenerate electrons, potentially allowing for greater mass accumulation Kalita et al. (2021a, b). Notably, super-Chandrasekhar mass WDs have not been directly observed in surveys such as Gaia or Kepler, likely because of their expected low luminosities.
The existence of over-luminous SNe Ia raises questions regarding the reliability of SNe Ia as standard candles, given their exceptionally high luminosities. Moreover, these events exhibit faster light curve decay rates compared to the standard SNe Ia, deviating from the standard Phillips relation Taubenberger (2017). As previously mentioned, the theoretical understanding of the explosion mechanisms of these over-luminous events remains incomplete, and there is no definitive observational evidence to accurately determine the progenitor mechanism. In this study, we utilize a sample of eight over-luminous SNe Ia to calculate $H_{0}$ independent of any underlying cosmological model. We demonstrate that the inclusion of these over-luminous SNe Ia in the analysis yields a lower value for $H_{0}$, aligning more closely with measurements from early-universe observations.
In general, existence of over-luminous SNe Ia leads to their distances being underestimated. To correct for this, one needs to reduce the value of $H_{0}$ from the one obtained from the local measurement, making it more agreeable to the value obtained from CMB. Processes leading to progenitors exceeding the Chandrasekhar mass limit may lead to new dynamics within the progenitor and thereby to the synthesis of new elements. The exact detail of this is yet unknown and needs further exploration. This article is structured as follows. In Section 2, we review the fundamental properties of SNe Ia at peak luminosity and examine the dependence of their peak brightness on the synthesized nickel mass. Section 3 presents our dataset of over-luminous SNe Ia, and details the methodology employed to estimate $H_{0}$. Additionally, we perform a Bayesian analysis incorporating various priors to infer $H_{0}$ in a cosmology-independent manner. Finally, in Section 4, we discuss our findings and provide concluding remarks.
2 Revisiting luminosity of type Ia supernova
SNe Ia are used as one of the standard candles in astronomy to measure luminosity distances due to their predicting behavior in peak brightness and declining rates. If $L_{\text{SN\,Ia}}$ is the luminosity of the SN Ia, $F$ is its flux measured on the Earth, and $d$ is its luminosity distance, they are related as
$F=\frac{L_{\text{SN\,Ia}}}{4\pi d^{2}}.$ (2.1)
Here $d$ can be estimated using the distance modulus formula given by Choudhuri (2010)
$\mu=\mathsf{m}-\mathsf{M}=5\log_{10}\left(\frac{d}{10\rm\,pc}\right),$ (2.2)
where $\mathsf{m}$ is the apparent magnitude and $\mathsf{M}$ is absolute magnitude of the SN Ia with $d$ measured in parsecs. The peak luminosity of SN Ia mostly depends on the amount of nickel-56 produced, whereas the afterglow is largely due to the radioactive decay of nickel to Cobalt to Iron, as follows:
[…]
The relation between the peak luminosity and the mass of nickel-56 produced during the process is given by Woosley et al. (2007)
$L_{\text{max}}\sim fM_{\text{Ni}}\exp\left(-\frac{t_{p}}{t_{\text{Ni}}}\right),$ (2.4)
[…]
$t_{d}\sim\Phi_{\text{Ni}}\kappa^{1/2}M_{\text{Ni}}^{3/4}E_{k}^{-1/4},$ (2.5)
[…]
$L_{\text{max}}=\left(6.45e^{\frac{-t_{r}}{8.8\,\mathrm{d}}}+1.45e^{\frac{-t_{r}}{111.3\,\mathrm{d}}}\right)\times 10^{43}\frac{M_{\text{Ni}}}{M_{\odot}}\rm\,erg\,s^{-1},$ (2.6)
where $t_{r}$ is the rising time of bolometric luminosity in days. Following Stritzinger et al. (2006), if we assume 19 days rising time in the bolometric luminosity with uncertainty of 3 days, the above equation reduces to
$L_{\text{max}}=(2.0\pm 0.3)\times 10^{43}\frac{M_{\text{Ni}}}{M_{\odot}}\rm\,erg\,s^{-1}.$ (2.7)
3 Hubble constant estimation using over-luminous type Ia supernovae
In this study, we analyze eight over-luminous SNe Ia, whose progenitors are thought to be super-Chandrasekhar WDs, and thereby estimate $H_{0}$. The relevant data for these eight supernovae are presented in Table 1. Here $z$ represents the redshift, $\mathsf{m}_{B}$ is the apparent bolometric magnitude, and $\mathsf{M}_{B}$ is the absolute bolometric magnitude at the peak of the SNe Ia. $M_{\mathrm{WD}}$ is the mass of the progenitor WD producing these SNe Ia inferred from the measured nickel mass and ejecta velocity. Note that to produce the measured value of the high nickel mass during the supernova explosion from a WD, its mass must exceed the Chandrasekhar mass limit.
Table 1: Data of over-luminous SNe Ia along with the estimated values of nickel mass and progenitor mass.
Name $z$ $\mathsf{M}_{B}$ $\mathsf{m}_{B}$ $M_{\mathrm{WD}}$ $M_{\mathrm{Ni}}$ Ref.
($M_{\odot}$) ($M_{\odot}$)
SN 2003fg 0.2440 -19.87 $\pm$ 0.06 20.35 $\pm$ 2.05* 2.1 1.3 Howell et al. (2006)
SN 2006gz 0.0237 -19.91 $\pm$ 0.21 16.06 $\pm$ 1.606* 2.0 1.2 Hicken et al. (2007)
SN 2007if 0.0742 -20.23 $\pm$ 2.023* 17.34 $\pm$ 0.04 2.4 1.5 Scalzo et al. (2010); Scalzo and Supernova Factory (2012)
SN 2009dc 0.0214 -20.22 $\pm$ 0.30 15.19 $\pm$ 0.16 2.4 1.4 Yamanaka et al. (2009); Silverman et al. (2011); Taubenberger et al. (2011)
SN 2012dn 0.010187 -19.52 $\pm$ 0.15 14.38 $\pm$ 0.02 1.6 0.82 Parrent et al. (2016); Chakradhari et al. (2014)
SN 2013cv 0.035 -19.84 $\pm$ 0.06 16.28 $\pm$ 0.03 1.59 0.81 Cao et al. (2016)
SN 2020esm 0.03619 -19.91 $\pm$ 0.15 16.16 $\pm$ 0.03 1.75 1.23 Dimitriadis et al. (2022)
LSQ 14fmg 0.0649 -19.87 $\pm$ 0.03 17.385 $\pm$ 0.008 1.45 1.07 Hsiao et al. (2020)
  • *Errors are taken to be 10% of the original value as their exact values are not reported.
  • **SN 2004gu is an over-luminous SN Ia as reported in Sun and Gal-Yam (2017). However, it was not included in the analysis due to the limited availability of observational data of the event.
3.1 Bayesian analysis to estimate the Hubble constant
Tripp provided a relation for calculating $H_{0}$ using $\mathsf{m}_{B}$ and $\mathsf{M}_{B}$, given in Tripp (1998)
$\log H_{0}$ $=\frac{\mathsf{M}_{B}-\mathsf{m}_{B}+5\log(\rm c\times 10^{6})}{5}+\log\left(\frac{1-q_{0}+q_{0}z-(1-q_{0})\sqrt{1+2q_{0}z}}{q_{0}^{2}}\right),$ (3.1)
where $q_{0}$ is the deceleration parameter and $c$ is the speed of light in $\rm km\,s^{-1}$ unit. Here both $z$ and $\mathsf{m}_{B}$ are observed quantities, whereas $\mathsf{M}_{B}$ is an inferred quantity, whose values are reported in Table 1 using the luminosity–distance relation with the assumption of $H_{0}$. Previous literature used $H_{0}$ in the range of $70-73\rm\,km\,s^{-1}\,Mpc^{-1}$ along with the assumption of a flat universe with values $q=-0.55$, $\Omega_{m}=0.3$, and $\Omega_{\Lambda}=0.7$ for the estimation of distance modulus. The peak absolute magnitude is related to the peak luminosity of Equation (2.7) and then plugging it into the luminosity–absolute magnitude relation as follows Carroll and Ostlie (1996)
$\mathsf{M}_{B}=\mathsf{M}_{\odot}-2.5\log\left(\frac{L_{\text{SN\,Ia}}}{L_{\odot}}\right).$ (3.2)
Moreover, using Equation (2.7), the amount of nickel produced during the event can be estimated considering the radioactive decay of nickel is the primary source driving the bolometric light curve. Although this holds true for the standard luminosity SNe Ia, new elements are thought to be synthesized in the over-luminous SNe Ia as the masses of the progenitors exceed the Chandrasekhar limit. It is expected that the luminosity of these over-luminous events may be affected by the other elements synthesized during the explosion. The exact connection between the luminosity and the mass of the progenitor is still unknown and hence the luminosity might not be captured completely by the radioactive decay of nickel alone. Thus, the aforementioned Tripp relation can be simplified as
$\log H_{0}$ $=\frac{\mathsf{M}_{B}-\mathsf{m}_{B}+52.38}{5}+\log\left(\frac{1.55-0.55z-1.55\sqrt{1-1.1z}}{0.3025}\right).$ (3.3)
According to this formula, $H_{0}$ primarily depends on three variables: $z$ and $\mathsf{m}_{B}$, which are observed quantities, while $\mathsf{M}_{B}$ is an inferred quantity. The absolute magnitude is usually estimated from the luminosities of regular SNe Ia using the Phillips relation, as shown in Phillips (1993). In the case of over-luminous SNe Ia, this relation becomes inadequate , as it tends to underestimate the intrinsic brightness of these events. Previous studies reported $\mathsf{M}_{B}$ values by assuming a fixed $H_{0}$ and employing the conventional luminsoity–distance relation. The Tripp formula estimates distance by incorporating the null geodesic distance derived from the the FLRW metric. Our goal here is to evaluate the absolute magnitude inferred from the Tripp formula and systematically compare it with the values reported in earlier literature. To reconcile discrepancies between the Tripp-derived $\mathsf{M}_{B}$ and previously reported estimates, we employ a Bayesian statistical framework to infer the optimal value of $H_{0}$. This method enables us to derive a refined and self-consistent estimate of $H_{0}$ that minimizes the variance in absolute magnitude assessments for over-luminous SNe Ia. Note that the sources are well-localized, leading to minimal errors in their redshifts.
Assuming Gaussian scatter of in $\mathsf{M}_{B}$ values, we now define the probability distribution for each SN Ia as
$P_{i}(\mathsf{m}_{B}\mid H_{0})=\frac{1}{\sqrt{2\pi\sigma_{i}^{2}}}\exp\left(\frac{\mathsf{M}_{B_{i}}-\mathsf{M}_{T_{i}}}{\sigma_{i}}\right)^{2},$ (3.4)
where $\sigma_{i}$ is the error bar for $\mathsf{M}_{B}$ and $\mathsf{M}_{T}$ is the expected absolute magnitude calculated by inverting Equation (3.3) and by assuming different values for the Hubble constant. Assuming a flat prior on $H_{0}$, the joint likelihood function can be defined as the product of each of the aforementioned individual likelihood function, given by
$\mathcal{L}$ $=\prod_{i=1}^{N}P_{i}(\mathsf{M}_{B}\mid H_{0}),$ (3.5)
[…]
$\mathcal{P}(H_{0})=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp{-\frac{(H_{0}-\bar{\mu})^{2}}{2\sigma^{2}}}.$ ) . (3.6)
We consider the mean $\bar{\mu}=73.0\rm\,km\,s^{-1}\,Mpc^{-1}$ and standard deviation $\sigma=1.4\rm\,km\,s^{-1}\,Mpc^{-1}$ following Freedman (2021) for Cepheid variable stars. Therefore, the joint likelihood function can be defined as
$\mathcal{L}=\prod_{i=1}^{8}\frac{1}{\sqrt{2\pi\sigma_{i}^{2}}}\exp\left(\frac{\mathsf{M}_{B_{i}}-\mathsf{M}_{T_{i}}}{\sigma_{i}}\right)^{2}\mathcal{P}(H_{0}).$ (3.7)
[…]
Table 2: Old and corrected distance modulus for over-luminous SNe Ia. Old $\mu$ are the values mentioned in the literature. $H_{0}$ is in the unit $\rm km\,s^{-1}\,Mpc^{-1}$. We consider flat $\Lambda$CDM universe with $\Omega_{m}=0.3$ and $\Omega_{\Lambda}=0.7$.
Name Old $\mu$ $H_{0}$ used Corrected $\mu$ Corrected $\mu$
($H_{0}=71.518$) ($H_{0}=68.100$)
SN2003fg 40.35 73.0 40.40 40.50
SN2006gz 34.98 73.0 35.02 35.13
SN2007if 37.60 71.0 37.58 37.69
SN2009dc 34.85 70.0 34.80 34.91
SN2012dn 33.16 72.0 33.17 33.28
SN2013cv 35.87 72.0 35.89 36.00
SN2020esm 35.92 73.0 35.96 36.07
LSQ14fmg 37.26 72.0 37.28 37.38
4 Discussion
This article discusses over-luminous SNe Ia and their potential use in understanding and resolving the Hubble tension. The use of over-luminous SNe Ia as standard candles affects the calibration of the distance ladder, and hence, can change the value of $H_{0}$. Using over-luminous SNe Ia leads to underestimating distances, and it is expected to bring the value of $H_{0}$ from $73.04\pm 1.04\rm\,km\,s^{-1}\,Mpc^{-1}$ as measured in the SH0ES survey, near that of the early-universe measurement. This study analyzes eight over-luminous SNe Ia listed in Table 1 to estimate $\mathsf{M_{B}}$ reported in previous studies with the same estimated using the Tripp formula. The goal is to understand the difference in the $\mathsf{M_{B}}$ values and determine $H_{0}$ for which the difference is minimized.
Utilizing a Bayesian statistical framework and a sample of eight over-luminous SNe Ia and eight normal luminosity SNe Ia, we demonstrate a reduction in $H_{0}$ from $73.04\pm 1.04\rm\,km\,s^{-1}\,Mpc^{-1}$ to $H_{0}=68.100\pm 1.048\rm\,km\,s^{-1}\,Mpc^{-1}$ with a flat prior and $H_{0}=71.518\pm 0.604\rm\,km\,s^{-1}\,Mpc^{-1}$ with the Gaussian prior based on Cepheid variable star measurements to estimate the distances. Moreover, combining them with standard luminosity SNe Ia, we achieve $H_{0}=70.622\pm 1.087\rm\,km\,s^{-1}\,Mpc^{-1}$ when using a flat prior. The Bayesian analysis yields significant deviation in $H_{0}$ estimate using over-luminous SNe Ia alone with repsct to the value reported by SH0ES collaboration, while a modest reduction in $H_{0}$ is obtained combining both over-luminous and standard one. This suggests that over-luminous SNe Ia can be reliably incorporated alongside standard SNe Ia in cosmological analyses for distance estimation. Nonetheless, the observed deviation underscores the need to extend the Phillips relation to account for over-luminous SNe Ia explicitly.
These results point toward the possibility that modifications to the standard FLRW metric may be necessary to resolve the persistent tension in the measurement of $H_{0}$. A detailed discussion of such a modification is provided in Krishnan et al. (2021), which demonstrates a reduction in $H_{0}$ to $71\rm\,km\,s^{-1}\,Mpc^{-1}$, a value that closely aligns with our Bayesian result based solely on over-luminous SNe Ia. Table 2 shows an increase in the distance modulus with a decrease in the value of $H_{0}$. This leads to the conclusion that the distances measured using the standard luminosity SNe Ia are an underestimation, and corrections are required to obtain the correct distances. However, this analysis introduces a novel discrepancy: the significant divergence between $H_{0}$ values obtained from over-luminous SNe Ia and those derived from regular SNe Ia. This raises fundamental questions regarding the intrinsic standardization of SNe Ia luminosities, particularly considering the inclusion of over-luminous events. The application of over-luminous SNe Ia as distance indicators may lead to a systematic underestimation of cosmological distances, contributing to their exclusion from standard candle applications. Furthermore, over-luminous SNe Ia exhibit substantial variability in peak luminosity, spectral characteristics, and light curve morphology. The potential existence of super-Chandrasekhar mass WD progenitors necessitates a rigorous re-evaluation of standardization methodologies for over-luminous SNe Ia, analogous to those applied to standard events. A comprehensive theoretical understanding of the explosion mechanisms associated with super-Chandrasekhar mass WDs, as suggested in Das and Mukhopadhyay (2013a), may facilitate the development of over-luminous SNe Ia as a distinct class of standardizable candles. A few attempts have been made to explain the WD to SN Ia mechanism in the single degenerate (WD accreting matter from a second star) case in Chen and Li (2009); Hachisu et al. (2012). A primary limitation of this study is the small sample size, which was constrained by the lack of available data. Future investigations employing an expanded catalog of over-luminous SNe Ia have the potential to refine the $H_{0}$ estimate and contribute substantially to the resolution of Hubble tension.
Acknowledgments
T.R.G. gratefully acknowledges Dr. Ravi Sheth from the University of Pennsylvania and Dr. S. Kalyana Rama from IMSc for insightful discussions and valuable comments. S.K. acknowledges funding by the National Center for Science, Poland, grant no. 2023/49/B/ST9/02777. A.R. and S.K. would like to thank IMSc, Chennai for providing support and resources during the course of this study.